OpenStudy (anonymous):

What is the x-intercept for this linear equation? 1/2x-2/3y=-6

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

Thx(:

jimthompson5910 (jim_thompson5910):

No it's not 1/2

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x 1/2x-2/3y=-6 1/2x-2/3(0)=-6 1/2x = -6 x = -6*2 x = -12 So the x-intercept is (-12,0)

OpenStudy (apoorvk):

X-intercept is the value of x for the equation when 'y'=0.